Consider the following demand 

Q=6000−3P


Find the maximum total revenue that the producer can get.


1
Expert's answer
2021-04-19T18:49:54-0400
"TR=pQ"

"p=2000-\frac{1}{3}Q"

"TR=2000Q-\frac{1}{3}Q^2"

"MR=2000-\frac{2}{3}Q"

"Q=3000"

"p=1000"

"TR=3000000"